OpenStudy (anonymous):

can u explain how 2 find x in 14/9=x

OpenStudy (amistre64):

whats wrong with x = 14/9 ???

OpenStudy (anonymous):

is this the equation?? \[\frac{14}{9} = x\]

OpenStudy (anonymous):

the equation is 14/9 = 6/x

OpenStudy (anonymous):

sry i thought i put the whole eqation up there

OpenStudy (amistre64):

multiply everyting by 9x/14

OpenStudy (anonymous):

\[\frac{14}{9} = \frac{6}{x}\] this is the same as this: \[\frac{9}{14} = \frac{x}{6}\] now multiply both sides by 6

OpenStudy (anonymous):

so u do 14x6 and 9x6 ?

OpenStudy (anonymous):

no, just the top line \[\frac{9 \times 6}{14} = x\]

OpenStudy (anonymous):

when i first learnt to manipulate equations i remember thinking about it visually: